OpenStudy (anonymous):

Solve the following system. x − y = 10 x + y = 8

sam (.sam.):

Elimination x − y = 10 x + y = 8 ----------------- -2y=2 y=-1 Substitute y=-1 into x + y = 8. x + y = 8 x + (-1) = 8 x-1=8 x=?

OpenStudy (anonymous):

x =9?

sam (.sam.):

yes , correct :)

sam (.sam.):

yes

OpenStudy (anonymous):

thanks :D

sam (.sam.):

welcome :)